# On Accessibility Fairness in Intermodal Autonomous Mobility-on-Demand Systems

### v2

@@ 2.1 Intermodal Network Structure
Similar to Salazar et al. (2020); Wollenstein-Betech et al. (2021), we model the mobility system and its different transportation modes on a digraph $\mathcal{G}=(\mathcal{V},\mathcal{A})$ consisting of a set of nodes $\mathcal{V}$ and a set of arcs $\mathcal{A}\subseteq\mathcal{V}\times\mathcal{V}$, which contain a car network layer $\mathcal{G}_{\mathrm{C}}=(\mathcal{V}_{\mathrm{C}},\mathcal{A}_{\mathrm{C}})$, a public transportation layer $\mathcal{G}_{\mathrm{P}}=(\mathcal{V}_{\mathrm{P}},\mathcal{A}_{\mathrm{P}})$, a bicycle layer $\mathcal{G}_{\mathrm{B}}=(\mathcal{V}_{\mathrm{B}},\mathcal{A}_{\mathrm{B}})$, a walking layer $\mathcal{G}_{\mathrm{W}}=(\mathcal{V}_{\mathrm{W}},\mathcal{A}_{\mathrm{W}})$, an origin layer $\mathcal{G}_{\mathrm{O}}=(\mathcal{V}_{\mathrm{O}},\emptyset)$ and a destination layer $\mathcal{G}_{\mathrm{D}}=(\mathcal{V}_{\mathrm{D}},\emptyset)$, which contain all origin and destination nodes, respectively, with no internal arcs. The car network consists of intersections $i\in\mathcal{V}_{\mathrm{C}}$ and car-driven road links $(i,j)\in\mathcal{A}_{\mathrm{C}}$. We model public transportation, i.e., tram and subway lines as distinct trees, using a set of station nodes $i\in\mathcal{V}_{\mathrm{P}}$ and a set of line segments $(i,j)\in\mathcal{A}_{\mathrm{P}}$. The bicycles and walking network consist of bike paths and walkable streets $(i,j)\in\mathcal{A}_{\mathrm{B}}$ and $(i,j)\in\mathcal{A}_{\mathrm{W}}$ between their intersections $i\in\mathcal{V}_{\mathrm{B}}$ and $i\in\mathcal{V}_{\mathrm{W}}$, respectively. Finally, we model the possibility of customers switching transportation modes (e.g., exiting the subway or hailing an AMoD ride) by connecting the origin layer to pedestrian, car, and bicycles layer, walking and bike layer to public transportation layer, and car, walking and biking layer to the destination layers. We capture all interconnections with a set of mode-switching arcs $\mathcal{A}_{\mathrm{S}}\subseteq\mathcal{V}_{\mathrm{C}}\times\mathcal{V}_{\mathrm{D}}\cup\mathcal{V}_{\mathrm{O}}\times\mathcal{V}_{\mathrm{C}}\cup\mathcal{V}_{\mathrm{P}}\times\mathcal{V}_{\mathrm{W}}\cup\mathcal{V}_{\mathrm{W}}\times\mathcal{V}_{\mathrm{P}}\cup\mathcal{V}_{\mathrm{B}}\times\mathcal{V}_{\mathrm{W}}\cup\mathcal{V}_{\mathrm{W}}\times\mathcal{V}_{\mathrm{B}}\cup\mathcal{V}_{\mathrm{O}}\times\mathcal{V}_{\mathrm{W}}\cup\mathcal{V}_{\mathrm{W}}\times\mathcal{V}_{\mathrm{D}}\cup\mathcal{V}_{\mathrm{O}}\times\mathcal{V}_{\mathrm{B}}\cup\mathcal{V}_{\mathrm{B}}\times\mathcal{V}_{\mathrm{D}}$. Additionally, layers do not share nodes with each other. Accordingly, $\mathcal{V}=\mathcal{V}_{\mathrm{W}}\cup\mathcal{V}_{\mathrm{B}}\cup\mathcal{V}_{\mathrm{C}}\cup\mathcal{V}_{\mathrm{P}}\cup\mathcal{V}_{\mathrm{O}}\cup\mathcal{V}_{\mathrm{D}}$ and $\mathcal{A}=\mathcal{A}_{\mathrm{W}}\cup\mathcal{A}_{\mathrm{B}}\cup\mathcal{A}_{\mathrm{C}}\cup\mathcal{A}_{\mathrm{P}}\cup\mathcal{A}_{\mathrm{S}}$ hold.
@@
Assuming no endogenous congestion effects, we consider a constant travel time vector $t_{a}\geq 0$ representing the time to traverse arc $a\in\mathcal{A}$, which represents the time to move (e.g., by car, bike, public transit, or walking) or to change mode of transportation (e.g., parking a bike, exiting an AMoD ride, waiting, getting on a public transit vehicle, etc.).
@@ 2.2 Operational Constraints and Time-optimal Problem
$\alpha_{m}\mathbbm{1}_{j=o_{m}}+\sum_{i:(i,j)\in\mathcal{A}}x^{m}_{(i,j)}=$ $\alpha_{m}\mathbbm{1}_{j=d_{m}}+\sum_{k:(j,k)\in\mathcal{A}}x^{m}_{(j,k)}$ (1)
$\forall m\in\mathcal{M}_{r},\forall r\in\mathcal{R},$
where $\mathbbm{1}$ is a Boolean indicator function defined as $\mathbbm{1}_{b}=1$ if $b$ is true and 0 otherwise. Similarly, the car balancing constraint is
@@
$=\sum_{k:(j,k)\in\mathcal{A}_{\mathrm{C}}}\left(x^{\mathrm{R}}_{(j,k)}+\sum_{r\in\mathcal{R}}\sum_{m\in\mathcal{M}_{r}}x^{m}_{(j,k)}\right).$
@@ 2.3 Accessibility Fairness Definition
which can be losslessly relaxed to convex as
@@
since it will be subject to minimization (Borsboom et al., 2021). We observe that such a definition captures the accessibility level of an o-d-pair on average. Specifically, if a fully accessible flow vector $x^{m}$ consists of more than one path, it could happen that some of those paths have a travel time above the threshold, whilst other paths have one below it, so that the weighted average travel time computed in (6) is below the threshold. Second, we define the accessibility unfairness level $u_{r}$ of a region $r$ as
@@
Problem 3 is an LP that can be efficiently solved with global optimality guarantees with off-the-shelf solution algorithms.
@@ 2.4 Discussion
A few comments are in order. First, we assume a constant travel time and no endogenous congestion effects. Such an assumption is in order for non car-based modes, and for relatively small fleets, where it is still possible to capture exogenous congestion effects (Rossi, 2018). Whilst the focus of the present paper is not on car-centric mobility, endogenous-congestion-aware traffic models could be included in future extensions (Salazar et al., 2019; Wollenstein-Betech et al., 2021). Second, we quantitatively capture accessibility unfairness in (7) using the flow-weighted extra time experienced by each region to accomplish its relevant trips. As accessibility is in general a qualitative multi-faceted concept from social science (Martens, 2017), we plan to study and possibly extend its quantitative definition to capture other aspects in future works. Finally, the slack variable definition in (6) captures the unaccessibility level of an o-d-pair on average, which corresponds to the path-based accessibility if there is only one path per flow-variable, which may not always be the case due to the presence of constraint (4). In this paper, we optimize the accessibility of paths a posteriori with the maximum-accessibility path allocation algorithm detailed in Section 2.5 below.
@@ 3 Results
In this section, we study the performance achievable by an I-AMoD system when operated in an accessibility-aware fashion and compare it to a more standard time-optimal operation. To this end, we compare the results obtained by solving Problem 2 and Problem 1, i.e., minimum-accessibility-unfairness vs minimum-time operation, respectively, where we allocate optimal flows to paths via Problem 3, using Algorithm 1. We set our case-study in the city of Eindhoven, NL. For the road network we gather information from OpenStreetMaps (Haklay and Weber, 2008), whilst for the public transportion network we use GTFS (2019). We generate realistic demand data with ALBATROSS (A Learning BAsed Transportation Oriented Simulation System), developed in Arentze and Timmermans (2004); Rasouli et al. (2018), which creates o-d-demands at PC4 level, i.e., the first 4 digits of the Dutch postal code. Specifically, it provides potential destinations and the ratio of the population that would travel there. Multiplying such ratios with the population size of each PC4 region $r$ and the average number of trips per person during a day (CBS, 2024), we obtain the average number of daily trips from an origin PC4 to a destination as flows $\alpha_{m}$ for $m\in\mathcal{M}_{r}$ and each region $r\in\mathcal{R}$. Last, we relate every o-d-pair at PC4 level to node level by evenly spreading the o-d-pairs to every node inside the origin and destination PC4. After mild clustering (see Appendix in Paparella et al. (2024a)), the resulting intermodal network has about 1300 nodes and 5000 arcs. The parameters used in the case-study are described in Table 1, where we select a fleet of $N_{\mathrm{cars,max}}=4000\,\mathrm{cars}$, and an accessibility threshold of $T_{\mathrm{max}}=20\,\mathrm{min}$.
@@
$t_{a}$, $a\in\mathcal{V}_{\mathrm{W}}\times\mathcal{V}_{\mathrm{P}}$ from GTFS (2019) min
@@ 3.1 Performance of Minimum-accessibility-unfairness vs Minimum-travel-time I-AMoD Operation
Travel Time [min] $12.34$ $12.39$
Unfairness Level per o-d-pair [min] $0.1483$ $0.0004$
Unfairness Level per path [min] $0.1610$ $0.1338$
Whilst the time-optimal operation achieves the lowest travel time, the minimum-accessibility-unfairness operation manages to shift the average travel time of each o-d-pair below the threshold $T_{\mathrm{max}}$, completely reaching accessibility fairness at the cost of a marginal increase in average travel time below 1%. The same result is shown in Fig. 6, depicting the average accessibility unfairness for the population of the PC4 regions $u_{r}$. Fig. 3 shows the difference between the two histograms, revealing that cars are subtracted at the cost of potentially longer rebalancing distances from already fast routes that would be accessible also by bike and other modes, and used to reduce accessibility unfairness of o-d-pairs that are further away. Fig. 7 shows this phenomenon for two exemplary o-d-pairs.
Refer to captionRefer to caption
@@ 4 Conclusion
We thank Dr. Ilse New for proofreading this paper.

### v3

@@ 2.1 Intermodal Network Structure
Similar to Salazar et al. (2020); Wollenstein-Betech et al. (2021), we model the mobility system and its different transportation modes on a digraph $\mathcal{G}=(\mathcal{V},\mathcal{A})$ consisting of a set of nodes $\mathcal{V}$ and a set of arcs $\mathcal{A}\subseteq\mathcal{V}\times\mathcal{V}$, which contain a car network layer $\mathcal{G}_{\mathrm{C}}=(\mathcal{V}_{\mathrm{C}},\mathcal{A}_{\mathrm{C}})$, a public transportation layer $\mathcal{G}_{\mathrm{P}}=(\mathcal{V}_{\mathrm{P}},\mathcal{A}_{\mathrm{P}})$, a bicycle layer $\mathcal{G}_{\mathrm{B}}=(\mathcal{V}_{\mathrm{B}},\mathcal{A}_{\mathrm{B}})$, a walking layer $\mathcal{G}_{\mathrm{W}}=(\mathcal{V}_{\mathrm{W}},\mathcal{A}_{\mathrm{W}})$, an origin layer $\mathcal{G}_{\mathrm{O}}=(\mathcal{V}_{\mathrm{O}},\emptyset)$ and a destination layer $\mathcal{G}_{\mathrm{D}}=(\mathcal{V}_{\mathrm{D}},\emptyset)$, which contain all origin and destination nodes, respectively, with no internal arcs. The car network consists of intersections $i\in\mathcal{V}_{\mathrm{C}}$ and car-driven road links $(i,j)\in\mathcal{A}_{\mathrm{C}}$. We model public transportation, i.e., tram and subway lines as distinct trees, using a set of station nodes $i\in\mathcal{V}_{\mathrm{P}}$ and a set of line segments $(i,j)\in\mathcal{A}_{\mathrm{P}}$. The bicycles and walking network consist of bike paths and walkable streets $(i,j)\in\mathcal{A}_{\mathrm{B}}$ and $(i,j)\in\mathcal{A}_{\mathrm{W}}$ between their intersections $i\in\mathcal{V}_{\mathrm{B}}$ and $i\in\mathcal{V}_{\mathrm{W}}$, respectively. Finally, we model the possibility of customers switching transportation modes (e.g., exiting public transport or hailing an AMoD ride) by connecting the origin layer to walking, car, and bicycles layer, walking to bike, public transportation and car layer, and car, walking and biking layer to the destination layers. We capture all interconnections with a set of mode-switching arcs $\mathcal{A}_{\mathrm{S}}\subseteq\mathcal{V}_{\mathrm{C}}\times\mathcal{V}_{\mathrm{D}}\cup\mathcal{V}_{\mathrm{O}}\times\mathcal{V}_{\mathrm{C}}\cup\mathcal{V}_{\mathrm{P}}\times\mathcal{V}_{\mathrm{W}}\cup\mathcal{V}_{\mathrm{W}}\times\mathcal{V}_{\mathrm{P}}\cup\mathcal{V}_{\mathrm{B}}\times\mathcal{V}_{\mathrm{W}}\cup\mathcal{V}_{\mathrm{O}}\times\mathcal{V}_{\mathrm{W}}\cup\mathcal{V}_{\mathrm{W}}\times\mathcal{V}_{\mathrm{D}}\cup\mathcal{V}_{\mathrm{O}}\times\mathcal{V}_{\mathrm{B}}\cup\mathcal{V}_{\mathrm{B}}\times\mathcal{V}_{\mathrm{D}}$. Additionally, layers do not share nodes with each other. Accordingly, $\mathcal{V}=\mathcal{V}_{\mathrm{W}}\cup\mathcal{V}_{\mathrm{B}}\cup\mathcal{V}_{\mathrm{C}}\cup\mathcal{V}_{\mathrm{P}}\cup\mathcal{V}_{\mathrm{O}}\cup\mathcal{V}_{\mathrm{D}}$ and $\mathcal{A}=\mathcal{A}_{\mathrm{W}}\cup\mathcal{A}_{\mathrm{B}}\cup\mathcal{A}_{\mathrm{C}}\cup\mathcal{A}_{\mathrm{P}}\cup\mathcal{A}_{\mathrm{S}}$ hold.
@@
Assuming no endogenous congestion effects, we consider a constant travel time vector $t_{a}\geq 0$ representing the time needed to traverse arc $a\in\mathcal{A}$, which captures the time to move (e.g., by car, bike, public transit, or walking) or to change mode of transportation (e.g., parking a bike, exiting an AMoD ride, waiting, getting on a public transit vehicle, etc.).
@@ 2.2 Operational Constraints and Time-optimal Problem
$\alpha_{m}\mathds{1}_{j=o_{m}}+\sum_{i:(i,j)\in\mathcal{A}}x^{m}_{(i,j)}=$ $\alpha_{m}\mathds{1}_{j=d_{m}}+\sum_{k:(j,k)\in\mathcal{A}}x^{m}_{(j,k)}$ (1)
$\forall j\in\mathcal{V},\forall m\in\mathcal{M}_{r},\forall r\in\mathcal{R},$
where $\mathds{1}$ is a Boolean indicator function defined as $\mathds{1}_{b}=1$ if $b$ is true and 0 otherwise. Similarly, the car balancing constraint is
@@
$=\sum_{k:(j,k)\in\mathcal{A}_{\mathrm{C}}}\left(x^{\mathrm{R}}_{(j,k)}+\sum_{r\in\mathcal{R}}\sum_{m\in\mathcal{M}_{r}}x^{m}_{(j,k)}\right)\;\forall j\in\mathcal{V}_{\mathrm{C}}.$
@@ 2.3 Accessibility Fairness Definition
which can be losslessly relaxed to a convex constraint as
@@
since it will be subject to minimization (Borsboom et al., 2021). We observe that such a definition captures the accessibility level of an o-d-pair on average. Specifically, if a fully accessible flow vector $x^{m}$ consists of more than one path, it could happen that some of those paths have a travel time above the threshold, whilst other paths have one below it, so that the weighted average travel time computed in (6) is below the threshold. Second, we define the accessibility unfairness level $u_{r}$ of a region $r$ as the average of extra travel time experienced for each o-d-pair belonging to it, weighted with the frequency of occurrence:
@@
Problem 2 is an LP that can be efficiently solved with global optimality guarantees with off-the-shelf solution algorithms.
@@ 2.4 Discussion
A few comments are in order. First, we assume a constant travel time and no endogenous congestion effects. Such an assumption is in order for non car-based modes, and for relatively small fleets, whereby it readily captures exogenous congestion effects (Rossi, 2018). Whilst the focus of the present paper is not on car-centric mobility, endogenous-congestion-aware traffic models could be included in future extensions (Salazar et al., 2019; Wollenstein-Betech et al., 2021). Second, we quantitatively capture accessibility unfairness in (7) using the flow-weighted extra time experienced by each region to accomplish its relevant trips. As accessibility is in general a qualitative multi-faceted concept from social science (Martens, 2017), we plan to study and possibly extend its quantitative definition to capture other aspects in future works. Finally, the slack variable definition in (6) captures the unaccessibility level of an o-d-pair on average, which corresponds to the path-based accessibility if there is only one path per flow-variable, which may not always be the case due to the presence of constraint (4). In this paper, we optimize the accessibility of paths a posteriori with the maximum-accessibility path allocation algorithm detailed in Section 2.5 below.
@@ 3 Results
In this section, we study the performance achievable by an I-AMoD system when operated in an accessibility-aware fashion and compare it to a more standard time-optimal operation. To this end, we compare the results obtained by solving Problem 2 and Problem 1, i.e., minimum-accessibility-unfairness vs minimum-time operation, respectively, where we allocate optimal flows to paths via Problem 3, using Algorithm 1. We set our case-study in the city of Eindhoven, NL. For the road network we gather information from OpenStreetMaps (Haklay and Weber, 2008), whilst for the public transportion network we use GTFS (2019). We generate realistic demand data with ALBATROSS (A Learning BAsed TRansportation Oriented Simulation System), developed in Arentze and Timmermans (2004); Rasouli et al. (2018), which creates o-d-demands at PC4 level, i.e., the first 4 digits of the Dutch postal code. Specifically, it provides potential destinations and the ratio of the population that would travel there. Multiplying such ratios with the population size $n_{r}$ of each PC4 region $r$ and the average number of trips per person during a day (CBS, 2024), we obtain the average number of daily trips from an origin PC4 to a destination as flows $\alpha_{m}$ for $m\in\mathcal{M}_{r}$ and each region $r\in\mathcal{R}$. Last, we relate every o-d-pair at PC4 level to node level by evenly spreading the o-d-pairs to every node inside the origin and destination PC4. After mild clustering (see Appendix in Paparella et al. (2024a)), the resulting intermodal network has about 1300 nodes and 5000 arcs. The parameters used in the case-study are described in Table 1, where we select a fleet of $N_{\mathrm{cars,max}}=4000\,\mathrm{cars}$, and an accessibility threshold of $T_{\mathrm{max}}=20\,\mathrm{min}$.
@@
$t_{a}$, $a\in\mathcal{V}_{\mathrm{W}}\times\mathcal{V}_{\mathrm{P}}$ $8$ min
@@ 3.1 Performance of Minimum-accessibility-unfairness vs Minimum-travel-time I-AMoD Operation
Travel Time [min] $12.51$ $12.62$
Unfairness Level per o-d-pair [min] $0.2274$ $0.0645$
Unfairness Level per path [min] $0.2506$ $0.1947$
Whilst the time-optimal operation achieves the lowest travel time, the minimum-accessibility-unfairness operation manages to shift the average travel time of each o-d-pair below the threshold $T_{\mathrm{max}}$, almost completely reaching accessibility fairness at the cost of a marginal increase in average travel time below 1%. The same result is shown in Fig. 6, depicting the average accessibility unfairness for the population of the PC4 regions $u_{r}$. Fig. 3 shows the difference between the two histograms, revealing that cars are subtracted at the cost of potentially longer rebalancing distances from already fast routes that would be accessible also by bike and other modes, and used to reduce accessibility unfairness of o-d-pairs that are further away. Fig. 7 shows this phenomenon for two exemplary o-d-pairs.
@@ 4 Conclusion
We thank Dr. Ilse New for proofreading this paper. Furthermore, we are grateful to Prof. Frauke Behrendt and Prof. Karel Martens for the fruitful and inspiring discussions.